the plasma density on lower ionosphere has been measured during satellite reentry to be about 10^18 m at 50km altitude,10^17 at 70km, and 10^14 at 18km. what are the plasma frequencies there?
1
Expert's answer
2020-07-01T15:42:21-0400

If we ignore thermal motion of electrons and assume that the effective mass of electrons equals to the rest mass of electron, the plasma frequency only depends on the concentration:

"f(n)=\frac{\omega(n)}{2\pi}=\frac{e}{2\pi}\sqrt{\frac{n}{m_e\epsilon_0}}."

In this equation, e is the charge of electron.

Thus:


"\begin{vmatrix}
 50\text{ km} & 8.979\text{ GHz}\\
 70\text{ km} & 2.839\text{ GHz}\\
8\text{ km}&89.79\text{ MHz}
\end{vmatrix}"